OpenStudy (anonymous):

BC is ___ feet long? a.14 b.25 c.7√3 d.14√3

OpenStudy (anonymous):



OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@mathstudent55 @ganeshie8 @amistre64

OpenStudy (anonymous):

You are working with a 30-60-90 degree right triangle which is one of the "special" triangles whose trig ratios you will eventually have to commit to memory. One of the ratios necessary here is derived from the tan function. tan 30 = [sqrt(3)] / 3 tan 30 = (AC)/(BC) -> (BC) = (AC)/(tan 30) So, make the substitutions for AC and tan 30 and then simplify.

OpenStudy (anonymous):

Hint: for tan 30, you might find it a little easier to use the equivalent: tan 30 = 1 / [sqrt(3)]

OpenStudy (anonymous):

25? I think I did it wrong.

OpenStudy (anonymous):

Well, let's see what you did and then I can see how to help you.

OpenStudy (anonymous):

Well I don't member what I did Just tried to put the nmbers in the calculator but it came out wrong Im pretty sure

OpenStudy (anonymous):

np, just type out the substitutions for: (BC) = (AC)/(tan 30) So, make the substitutions for AC and tan 30 and then simplify. You'll then have a written equation for BC equals something. Write it out here so I can see it.

OpenStudy (anonymous):

To start, what is AC ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Good, and since I already wrote out what tan 30 is, you can make the substitutions into the right-hand side of: (BC) = (AC)/(tan 30) Write out the right side with substitutions, but without simplifying. I'll then be able to see what you're doing right or wrong.